# TriNeRFLet: A Wavelet Based Multiscale Triplane NeRF Representation

### v1

@@ 6 Conclusions
This paper introduced a new NeRF structure that relies on multiscale wavelet representation. This structure improved the performance of Triplane, which lagged behind NeRF SOTA methods, achieving competitive results. Having the multiscale structure enabled us to implement a new diffusion-based super-resolution for NeRF.
Despite the advantages our proposed approach brings, it also has some limitations, mainly due to the additional runtime overhead it brings to training time. This can be partially mitigated by using its light-weight versions and using the coarse to fine training, which reduces training time.
Our proposed approach of learning the features in the wavelet domain instead of the original one, opens the door for many future works to implement it in other applications. We believe this approach can benefit tasks where data is sparse or training suffers from over-fitting, by regularizing the high frequencies in the wavelet domain.
Appendix A Ablation
For all the experiments we conducted in the paper, we applied the Biorthogonal 6.8 (Bior6.8) wavelet. We present here an ablation experiment that checks the impact of other wavelet filters. We compare the reconstruction performance of 4 scenes from the Blender dataset. The setups we tested are vanilla Triplane (no wavelet), Haar, Biorthogonal 2.2 (Bior2.2), Biorthogonal 2.6 (Bior2.6), Biorthogonal 4.4 (Bior4.4) and Biorthogonal 6.8 (Bior6.8). The results are presented in Figure 9. We use the TriNeRFLet Base Light setting. Vanilla Triplane has the same size and structure as this setting. Note that all wavelets achieve a significant performance advantage over vanilla Triplane (as shown also in the paper). Generally, higher order wavelets provide a better representation of smooth functions [24]. Thus, it is not surprising that Bior6.8, Bior4.4 and Bior 2.6 achieve better performance than Bior 2.2 and Haar. Note though that in terms of training time, training with Haar wavelet is faster by up to $30\%$ compared to the other wavelets as Haar complexity is $O(N)$ and the complexity of applying the other wavelets is $O(N\log(N)$.

### v2

@@ 6 Conclusions
This paper introduced a new NeRF structure that relies on a multiscale wavelet representation. This structure improved the performance of Triplane, which lagged behind NeRF state-of-the-art methods, achieving competitive results. Having the multiscale structure enabled us to implement a new diffusion-guided SR for NeRF. We believe that the concept of learning the features in the multiscale wavelet domain instead of the original one has the potential to improve other vision-related applications.
Despite the advantages of our approach, it also has some limitations, mainly due to the runtime overhead it adds to training time. This can be partially mitigated by using its light-weight versions and the coarse to fine training, which reduce training time. We defer further run-time improvements to future work.
Acknowledgments This work was supported in part by KLA grant.
Supplementary Materials
Appendix 0.A Ablation
Wavelet type. For all the experiments we conducted in the paper, we applied the Biorthogonal 6.8 (Bior6.8) wavelet. We present here an ablation experiment that checks the impact of other wavelet filters. We compare the reconstruction performance of 4 scenes from the Blender dataset. The setups we tested are vanilla Triplane (no wavelet), Haar, Biorthogonal 2.2 (Bior2.2), Biorthogonal 2.6 (Bior2.6), Biorthogonal 4.4 (Bior4.4) and Biorthogonal 6.8 (Bior6.8). The results are presented in Figure 7. We use the TriNeRFLet Base Light setting. Vanilla Triplane has the same size and structure as this setting. Note that all wavelets achieve a significant performance advantage over vanilla Triplane (as shown also in the paper). Generally, higher order wavelets provide a better representation of smooth functions [25]. Thus, it is not surprising that Bior6.8, Bior4.4 and Bior 2.6 achieve better performance than Bior 2.2 and Haar. Note though that in terms of training time, training with Haar wavelet is faster by up to $30\%$ compared to the other wavelets as Haar complexity is $O(N)$ and the complexity of applying the other wavelets is $O(N\log(N))$.
@@ Appendix 0.A Ablation
Coarse To Fine. As mentioned in the paper, coarse to fine (c2f) accelerates TriNeRFLet framework and is presented as a mitigation for the wavelet inverse transform overhead. To further check the additional contributions of c2f, we conduct the Blender reconstruction experiment (Base Light version) but without c2f. Results are presented in Table 5, and as noticed, c2f indeed brings a contribution in the reconstruction quality, alongside the contribution in training time.
Wavelet Regularization. To check the importance of the wavelet regularization loss, we extend the c2f experiment above to be without wavelet regularization. The results in Table 5 illustrate the important contribution of wavelet regularization and that it is an essential component of TriNeRFLet performance.
Appendix 0.B Technical Details
3D Reconstruction. Table 4 contains the parameters for the reconstruction versions.
Blender Super-Resolution. In both Blender settings ($100\rightarrow 400\And 200\rightarrow 800$), we use $C=16$ (channels), $s_{LR}=6000$ (LR only steps), $s=16000$ (total steps), $s_{refresh}=500$ (refresh steps), $T_{min}=0.02$ and $T_{max}$ is initialized with $0.98$ and decreases linearly to reach $0.25$ at the end. We use [12] as a base for implementing the super-resolution scheme, as it is more suitable for this task. We run the SR method on a single A6000600060006000 RTX GPU, and it runs for almost $7$ hours.
LLFF Super-Resolution. We use similar settings to Blender $200\rightarrow 800$ experiment except for, $C=32$ (channels), $s=20000$ (total steps), $s_{refresh}=250$ (refresh steps). Like [28], we also operate in NDC space coordinates, as it fits this type of scenes better. Similar to the Blender experiment, we ran LLFF SR method on a single A6000 RTX GPU, where each scene runs almost for $12$ hours.
$N_{LL}$ $L$ $N_{base}$ $N_{final}$ $C$ $\gamma$ $W$ $D_{dens}$ $D_{col}$ train steps trainable parameters
Small $64$ $4$ $512$ $1024$ $16$ $0.2$ $64$ $1$ $2$ $6k$ $17M$
Base Light $64$ $5$ $512$ $2048$ $32$ $0.4$ $64$ $1$ $2$ $10k$ $134M$
Base $64$ $5$ $512$ $2048$ $32$ $0.4$ $64$ $1$ $2$ $43k$ $134M$
Large $64$ $5$ $512$ $2048$ $48$ $0.6$ $128$ $1$ $2$ $83k$ $201M$
Table 4: The configurations used in the different TriNeRFLet versions
Mic Chair Ship Materials Lego Drums Ficus Hotdog Avg. Train Time
Base Light $35.77$ $35.00$ $31.10$ $29.35$ $36.44$ $25.98$ $33.96$ $36.93$ $33.07$ $1.5$ hours
Base Light w/o c2f $35.52$ $35.08$ $31.11$ $29.14$ $36.48$ $25.76$ $33.46$ $36.88$ $32.93$ $2$ hours
[…]
Triplane $33.85$ $32.83$ $29.58$ $28.15$ $34.70$ $24.86$ $30.35$ $35.80$ $31.26$ $2$ hours
Table 5: Coarse to fine (c2f) and wavelet regularization importance ablation. We compare Base Light version with the versions without c2f and without c2f and wavelet regularization. The results indicate the important contribution of the c2f and wavelet regularization.
Appendix 0.C Detailed Results
Tables 6,7 and 8 contain the detailed results of the SR experiments described in the paper.
Mic Chair Ship Materials Lego Drums Ficus Hotdog Avg.
PSNR $31.54$ $29.4$ $28.11$ $27.9$ $27.65$ $24.18$ $27.7$ $31.92$ $28.55$
LPIPS $0.05$ $0.051$ $0.118$ $0.048$ $0.073$ $0.067$ $0.044$ $0.039$ $0.061$
SSIM $0.914$ $0.946$ $0.844$ $0.921$ $0.914$ $0.896$ $0.936$ $0.939$ $0.913$
Table 6: Blender $100\rightarrow 400$ detailed results.
Mic Chair Ship Materials Lego Drums Ficus Hotdog Avg.
PSNR $31.11$ $29.86$ $28.14$ $28.02$ $30.46$ $24.17$ $30.6$ $33.58$ $29.49$
LPIPS $0.032$ $0.059$ $0.114$ $0.046$ $0.032$ $0.071$ $0.021$ $0.034$ $0.051$
SSIM $0.966$ $0.944$ $0.844$ $0.93$ $0.922$ $0.913$ $0.96$ $0.96$ $0.93$
Table 7: Blender $200\rightarrow 800$ detailed results.
Flower Fern Leaves Fortress Horns Orchids Room Trex Avg.
PSNR $26.6$ $24$ $20.01$ $28.88$ $24.93$ $21.38$ $28.56$ $25.57$ $25.00$
LPIPS $0.214$ $0.205$ $0.239$ $0.175$ $0.26$ $0.244$ $0.125$ $0.163$ $0.203$
SSIM $0.808$ $0.766$ $0.639$ $0.832$ $0.737$ $0.668$ $0.901$ $0.819$ $0.771$
Table 8: LLFF $378\times 504\rightarrow 1512\times 2016$ detailed results.
Appendix 0.D Qualitative Results
Figures 8 and 9 provide qualitative results for reconstruction and LLFF SR experiments respectively. For all results, the reader is referred to the project webpage https://rajaeekh.github.io/trinerflet-web. Furthermore, TriNeRFLet-SR can be incorporated into a generative model. To show that, we use the diffusion process in [17] that generates a triplane and then apply SR to it using our scheme. Fig. 10 presents examples of triplanes rendered with resolution $128$ that are upscaled to $512$ by TriNeRFLet SR.
Triplane TriNeRFLet INGP Ground-Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
LR NeRF-SR TriNeRFLet HR
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: NeRF LLFF super-resolution qualitative results.
Figure 10: SR of low-res objects generated by triplane diffusion with text “Astronaut suit and helmet" and “Colorful electric scooter”. First row is low-res and second row is TriNerFLet-SR.